# Multi-language Unit Test Generation using LLMs

### v1

@@ 7. Threats to Validity
Internal threat. One potential threat to internal validity is the limited number of evaluation runs. While previous studies have performed more than ten runs, the substantial cost associated with running evaluations across several large applications and six different LLMs led us to limit the runs. Another threat can be automated naturalness evaluation, and to mitigate we conducted a survey of professional developers, and we found that the findings are very similar.
Construct threat. To ensure construct validity, we measured coverage and mutation scores using widely used tools such as JaCoCo (EclEmma, [n. d.]), Coverage.py (Batchelder, [n. d.]), PIT (PITEST, [n. d.]), and Mutmut (Hovmöller, [n. d.]).
8. Summary and Future Work
@@
In this paper, we presented aster, a multi-lingual test-generation tool that leverages LLMs guided by lightweight static analysis to generate natural and effective unit test cases for Java and Python. Through its preprocessing component, aster ensures that LLM prompts have adequate context required for generating unit tests for a focal method. aster’s postprocessing component performs iterative test repair and coverage augmentation. Our extensive evaluation, with six LLMs on a dataset of Java SE, Java EE, and Python application, showed that aster is competitive with state-of-the-art tools in coverage achieved on Java SE application, and outperforms them significantly on Java EE and Python applications, while also producing considerably more natural tests than those tools. Our developer survey, with over 160 participants, highlighted the naturalness characteristics of aster-generated tests and their usability for building automated test suites. Future research directions include extending aster to other PLs and levels of testing (e.g., integration testing), creating fine-tuned models for testing to reduce the cost of LLM interactions, and exploring techniques for improving fault-detection ability of the generated tests.

### v3

@@ VII Threats to Validity
To address threat related to generalizability of aster, we extended support for (a) multiple PLs and (b) multiple models with varying sizes, modalities, and families. Another potential threat is the limited number of evaluation runs. While previous studies have performed $>10$ runs, the substantial cost associated with running evaluations across several applications and six LLMs led us to limit the runs. Another potential threat is the automated naturalness evaluation; to mitigate that, we conducted a survey of professional developers, and found that the findings are very similar.
VIII Summary and Future Works
@@
In this paper, we presented aster, a multi-language test-generation tool that leverages LLMs guided by lightweight static analysis to generate natural and effective unit test cases for Java and Python. Through its preprocessing component, aster ensures that LLM prompts have adequate context required for generating unit tests for a focal method. aster’s postprocessing component performs iterative test repair and coverage augmentation. Our extensive evaluation, with six LLMs on a dataset of Java SE, Java EE, and Python application, showed that aster is competitive with state-of-the-art tools in coverage achieved on Java SE application, and outperforms them significantly on Java EE and Python applications, while also producing considerably more natural tests than those tools. Our developer survey, with over 160 participants, highlighted the naturalness characteristics of aster-generated tests and their usability for building automated test suites. Future research directions include extending aster to other PLs and levels of testing (e.g., integration testing), creating fine-tuned models for testing to reduce the cost of LLM interactions, and exploring techniques for improving fault-detection ability of the generated tests.
Acknowledgment
We thank Xuan Liu, Maja Vuković, Nicholas Fuller, Angel Montesdeoca, Lisa Waugh, Jayant Talekar, and all the members of IBM watsonx Code Assistant team for their help with this research. We also thank the anonymous reviewers for their comments, which helped make this work stronger.